# Raman phonon dynamics and its control for enhanced optical frequency conversion

### 27 Mar 2026 v2

@@ 1 Introduction
Recently, we have developed a new perspective for understanding Raman dynamics based on the temporally-varying Raman-induced index modulation (Fig. 1) [13, 14], where is the Raman temporal response and is the optical field assuming an unchanged spatial profile, e.g., in an optical fiber. The nonlinear Raman field increment follows . Traditional approaches isolate only the oscillatory component of the index response by solving the coupled phonon equation with an imposed wave solution [: phonon wave vector; : phonon oscillation angular frequency which is equal to the Raman transition angular frequency] [15, 16, 17]. In contrast, this framework enables direct visualization of the full spectrotemporal Raman dynamics that combines frequency-dependent Raman gain spectrum and time-dependent phonon dynamics. In the long-pulse regime, the pulse-following component of contributes to Raman-enhanced self-phase modulation, facilitating supercontinuum generation and ultrashort-pulse formation [18, 19, 20, 21]. In the ultrashort-pulse regime, the pulse impulsively drives the Raman scattering whose inertial induces a significantly-delayed index modulation such that the pulse interacts with only the initial rising edge of the index wave. Operation in the intermediate regime combines features of both limits, yielding enhanced spectral broadening accompanied by a net redshift [22, 23, 24]. Moreover, the slow inertial response smooths the nonlinear index profile, facilitating high-fidelity spectral compression [25]. Although this framework provides a clear and comprehensive understanding of Raman dynamics across different temporal regimes, it leads to a new question regarding the connection between the index modulation and the traditional concept of quantized phonon. Is pulse-following index that also originates from Raman scattering a part of the Raman phonon? This article will explains the role of Raman phonon in terms of index modulation.
@@ 2 Raman phonon
with the reduced mass of a diatomic molecule, the vibrational transition angular frequency, the unperturbed population at level , the polarizability anisotropy, the real-valued electric field (with complex-valued representing its analytic signal), and the dephasing rate and the rotational transition angular frequency between rotational levels and , respectively. The coherence corresponds to the negative-frequency component of the displacement in vibrations, or directly proportional to the orientation in rotations, which, for example under linear polarization, becomes .
Phonons provide a quantum-mechanical description of the oscillatory component of Raman-induced index modulation (Fig. 3), namely only those oscillating around . In the impulsive regime, the induced index modulation exhibits a pronounced oscillatory behavior that persists beyond the duration of the driving pulse, a response that intuitively corresponds to the excitation of phonons. By contrast, in the long-pulse transient and steady-state regimes, the index predominantly follows the temporal envelope of the optical pulse and does not exhibit intrinsic oscillations. As pulse-following index does not induce an overall redshift for the pump but enhances spectrally-symmetric self-phase modulation, no net energy is transferred to the medium and thus no phonon is generated. Oscillatory behavior emerges only after Raman fields are generated, whose interference produces high-frequency beating that impulsively drives the oscillatory index response. The amplitude of these oscillations is determined by the strength of the optical beating, which in turn depends on the Stokes-field amplitude and its photon number. As one Stokes photon generated corresponds to one phonon generated, the magnitude of the oscillatory response reflects the number of phonons generated in the Raman process.
@@ 3 Linear phonon-mediated process
Distinct nonlinear Raman phenomena emerge when an optical field propagates in a medium containing a pre-existing population of phonons. In the impulsive regime with an ultrashort driving field, the interaction reduces to the familiar femtosecond pump-probe configuration: depending on the relative temporal delay, the pulse can be redshifted, blueshifted, or temporally compressed [55, 56, 57, 58, 59, 60]. In contrast, the dynamics in the long-pulse transient and steady-state regimes are governed by wave-vector matching between the existing phonons and the Raman interaction [31, 32, 13]. Because newly-generated (absorbed) phonons are inherently phase-matched to the Stokes (anti-Stokes) process that creates (annihilates) them, this condition can equivalently be interpreted as wave-vector matching between the existing phonons and the phonons to be generated (or absorbed). Such wave-vector matching plays a crucial role in multi-field Raman interactions, for example in two-pulse schemes employing different wavelengths [32], or in cascaded Stokes processes of different orders that generate phonons having different wave vectors. This process follows (Fig. 4; see Supplementary Sec. 10 in [13] for derivation details):
@@
This process also underlies the coherent Stokes generation from spontaneous Raman scattering in the transient regime, where the phonon lifetime is much longer than the pulse duration (Fig. 5) [76]. To investigate it, we simulated with the unidirectional pulse propagation equation (UPPE) that considers realistic Raman response function for several types of gases [13], and the result is shown in Fig. 5. Initially, the Stokes field is seeded by vacuum fluctuations [77]. Due to the incoherence of vacuum fluctuations, the beating with the pump field remains fluctuating, leading to incoherent phonon response. The Stokes increment during propagation [14] is therefore dominated by the pump field (top figure in Fig. 5). However, limited phonon response time (due to the Raman period) results in a smoothed phonon wave with increased coherence that further drives more coherent Stokes generation through the linear phonon-mediated process. This new coherent Stokes wave overrides the initial incoherent Stokes wave, seeding the latter nonlinear amplification and eventually producing a coherent Stokes field that inherits the temporal phase from the pump [Eq. (3a)]. Moreover, the growing phonon wave leads to a Stokes increment that is delayed to the trailing edge of the pump pulse. Traditional frameworks based on nonlinear Raman amplification, in which the incoherent Stokes background is amplified, cannot adequately account for the emergence of coherence in this process. et al.Carman analyzed transient Stokes generation by assuming a priori that the Stokes and pump fields share the same phase structure as they evolve in a high-gain environment [76]. Only with the introduction of the linear phonon-mediated dynamics is this coherent generation now clearly explained.
Eq. (3) is introduced as an abstraction layer that isolates the linear phonon-mediated contribution from the full Raman response , clarifying the specific structural role this pathway plays in Raman temporal dynamics. Because the phonon wave is continuously driven by the optical fields, its amplitude and phase are not fixed material parameters but evolve with propagation distance. This construction therefore does not constitute an independent evolution equation; the actual dynamics must still be obtained from the full Raman formulation where explicitly contains both the field components that excite the phonons and those that are to be converted. The newly-introduced equations thus complement, rather than supplant, the conventional nonlinear description by selectively filtering out the linear phonon-mediated pathway to make its physical role more transparent.

### 02 Apr 2026 v3

@@ 1 Introduction
Recently, we have developed a new perspective for understanding Raman dynamics based on the temporally-varying Raman-induced index modulation (Fig. 1) [13, 14], where is the Raman temporal response and is the optical field assuming an unchanged spatial profile, e.g., in an optical fiber. The nonlinear Raman field increment follows . Traditional approaches isolate only the oscillatory component of the index response by solving the coupled phonon equation with an imposed wave solution [: phonon wave vector; : phonon oscillation angular frequency which is equal to the Raman transition angular frequency] [15, 16, 17]. In contrast, this framework enables direct visualization of the full spectrotemporal Raman dynamics that combines frequency-dependent Raman gain spectrum and time-dependent phonon dynamics. In the long-pulse regime, the pulse-following component of contributes to Raman-enhanced self-phase modulation (SPM), facilitating supercontinuum generation and ultrashort-pulse formation [18, 19, 20, 21]. In the ultrashort-pulse regime, the pulse impulsively drives the Raman scattering whose inertial induces a significantly-delayed index modulation such that the pulse interacts with only the initial rising edge of the index wave. Operation in the intermediate regime combines features of both limits, yielding enhanced spectral broadening accompanied by a net redshift [22, 23, 24]. Moreover, the slow inertial response smooths the nonlinear index profile, facilitating high-fidelity spectral compression [25]. Although this framework provides a clear and comprehensive understanding of Raman dynamics across different temporal regimes, it leads to a new question regarding the connection between the index modulation and the traditional concept of quantized phonon. Is pulse-following index that also originates from Raman scattering a part of the Raman phonon? This article will explains the role of Raman phonon in terms of index modulation.
@@ 2 Raman phonon
with the reduced mass of a diatomic molecule, the vibrational transition angular frequency, the unperturbed population at level , the polarizability anisotropy, the real-valued electric field (with complex-valued representing its analytic signal [52]), and the dephasing rate and the rotational transition angular frequency between rotational levels and , respectively. The coherence corresponds to the negative-frequency component of the displacement in vibrations, or directly proportional to the orientation in rotations, which, for example under linear polarization, becomes .
Phonons provide a quantum-mechanical description of the oscillatory component of Raman-induced index modulation (Fig. 3), namely only those oscillating around which we denote as hereafter. In the impulsive regime, the induced index modulation exhibits a pronounced oscillatory behavior that persists beyond the duration of the driving pulse, a response that intuitively corresponds to the excitation of phonons. By contrast, in the long-pulse transient and steady-state regimes, the index predominantly follows the temporal envelope of the optical pulse and does not exhibit intrinsic oscillations. As pulse-following index does not induce an overall redshift for the pump but enhances spectrally-symmetric SPM, no net energy is transferred to the medium and thus no phonon is generated. Oscillatory behavior emerges only after Raman fields are generated, whose interference produces high-frequency beating that impulsively drives the oscillatory index response. The amplitude of these oscillations is determined by the strength of the optical beating, which in turn depends on the Stokes-field amplitude and its photon number. As one Stokes photon generated corresponds to one phonon generated, the magnitude of the oscillatory response reflects the number of phonons generated in the Raman process.
@@ 3 Linear phonon-mediated process
Distinct nonlinear Raman phenomena emerge when an optical field propagates in a medium containing a pre-existing population of phonons (Fig. 4). In the impulsive regime with an ultrashort driving field, the interaction reduces to the familiar femtosecond pump-probe configuration: depending on the relative temporal delay, the pulse can be redshifted, blueshifted, or temporally compressed [56, 57, 58, 59, 60, 61]. In contrast, the dynamics in the long-pulse transient and steady-state regimes are governed by wave-vector matching between the existing phonons and the Raman interaction [31, 32, 13]. Because newly-generated (absorbed) phonons are inherently phase-matched to the Stokes (anti-Stokes) process that creates (annihilates) them, this condition can equivalently be interpreted as wave-vector matching between the existing phonons and the phonons to be generated (or absorbed). Such wave-vector matching plays a crucial role in multi-field Raman interactions, for example in two-pulse schemes employing different wavelengths [32], or in cascaded Stokes processes of different orders that generate phonons having different wave vectors. This process follows (Fig. 5; see Supplementary Sec. 10 in [13] for derivation details):
@@
This process also underlies the coherent Stokes generation from spontaneous Raman scattering in the transient regime (Fig. 6) [77]. Unlike incoherent Stokes fields from steady-state spontaneous Raman generation, transient spontaneous Raman generation produces coherent Stokes fields that exhibit the same temporal phase structure as the pump pulse. To investigate it, we simulated with the unidirectional pulse propagation equation (UPPE) that considers realistic Raman response function for several types of gases [13], and the result is shown in Fig. 6. Initially, the Stokes field is seeded by vacuum fluctuations [78]. Due to the incoherence of vacuum fluctuations, the beating with the pump field remains fluctuating, leading to incoherent phonon response. The Stokes increment during propagation [14] is therefore dominated by the pump field (top figure in Fig. 6). However, limited phonon response time (due to the Raman period) results in a smoothed phonon wave with increased coherence that further drives more coherent Stokes generation through the linear phonon-mediated process. This new coherent Stokes wave overrides the initial incoherent Stokes wave, seeding the latter nonlinear amplification and eventually producing a coherent Stokes field that inherits the temporal phase from the pump [Eq. (3a)]. Moreover, the growing phonon wave leads to a Stokes increment that is delayed to the trailing edge of the pump pulse. Traditional frameworks based solely on nonlinear Raman amplification, in which the incoherent Stokes background is amplified, cannot adequately account for the emergence of coherence in this process. It omits the phonon temporal dynamics crucial for coherence generation and is thus applicable only in the steady-state regime. et al.Carman analyzed transient Stokes generation by assuming a priori that the Stokes and pump fields share the same phase structure as they evolve in a high-gain environment [77]. Only with the introduction of the linear phonon-mediated dynamics is this coherence generation now clearly explained.
Eq. (3) is introduced as an abstraction layer that isolates the linear phonon-mediated contribution from the full Raman response , clarifying the specific structural role this pathway plays in Raman temporal dynamics. Because the phonon wave is continuously driven by the optical fields, its amplitude and phase vary with propagation distance, whereas the governing phonon wave vector itself remains fixed as determined by the dispersion relation in the interacting medium. This construction therefore does not constitute an independent evolution equation; the actual dynamics must still be obtained from the full Raman formulation where explicitly contains both the field components that excite the phonons and those that are to be converted. The newly-introduced equations thus complement, rather than supplant, the conventional nonlinear description by selectively filtering out the linear phonon-mediated pathway to make its physical role more transparent.